Calculate the range of wavelengths that are received by the typical FM radio in a car
Answer #1

FM frequency range: 88 Mhz - 108 Mhz

Higher wavelength = c/f = 3 * 10^8 / (88 * 10^6) = 3.41 m

lower wavelength = c/f = 3 * 10^8 / (108 * 10^6) = 2.78 m

Thus, the range is [2.78, 3.41] m